# Exploiting Scale-Variant Attention for Segmenting Small Medical Objects

### v4

@@ IV-A1 Dataset
SpermHealth is a customized dataset from the Prince of Wales Hospital in Hong Kong consisting of low-resolution sperm images (640×480, 96 DPI) extracted from microscope-captured videos. These images have been meticulously annotated into normal and abnormal categories by experienced fertility doctors. Further details of the datasets used in the tests are presented in Tab. I.
@@ IV-D4 Inference Time
This section quantifies the inference characteristics of the tested networks, including the number of parameters (# Parameters), multiply-accumulate operations (MACs), and inferencing speed. The number of classes was set to eight, and other configurations were consistent with those described in Sec. IV-A2. The inference speed results, averaged over 1000 runs, are presented in Tab. VII. Tab. VII illustrates that SvANet achieved a real-time analysis of medical images with 46.91 frames per second (FPS), which is twice as fast as UNet, 46% faster than nnUNet, and 19% faster than TransNetR. This performance level is suitable for self-examination and clinical diagnostic applications.
# Parameters | MACs | Speed | |
Methods | /Million | /Billion | /FPS |
UNet (MICCAI’15) [10] | 34.53 | 200.75 | 22.76 |
UNet++ (TMI’19) [11] | 9.16 | 107.09 | 43.69 |
PraNet (MICCAI’20) [12] | 30.34 | 25.65 | 68.16 |
HRNet (TPAMI’20) [29] | 63.60 | 65.80 | 78.95 |
nnUNet (NM’21) [13] | 30.60 | 232.19 | 32.11 |
CFANet (PR’23) [16] | 25.71 | 115.63 | 60.89 |
TransNetR (MIDL’24) [25] | 27.27 | 44.71 | 39.39 |
SvANet (Ours) | 177.64 | 312.76 | 46.91 |
@@ IV-E Negative Case Studies
For tissue cell recognition in the TissueNet dataset, as shown in Fig. 9c, both TransNetR and SvANet effectively delineated cell boundaries and accurately labeled the cells and nuclei regions, closely resembling the ground truth. In contrast, other SOTA methods struggled to categorize cells and nuclei, leading to difficulties in differentiating cell boundaries and merging several cells into one. For sperm cell analysis, as presented by the final image in Fig. 9d, SvANet precisely located all sperm positions and effectively recognized the region of the broken-line tail of an abnormal sperm. Conversely, tested methods like HRNet struggled to distinguish the broken-line shape. Moreover, UNet, nnUNet, and TransNetR misclassified an unhealthy sperm head as healthy, as indicated by a green subregion in Fig. 9d.

### v5

@@ IV-A1 Dataset
SpermHealth is a customized dataset from the 3rd Affiliated Hospital of Shenzhen University, consisting of low-resolution sperm images (640×480, 96 DPI) extracted from microscope-captured videos. These images have been meticulously annotated into normal and abnormal categories by experienced fertility doctors. Further details of the datasets used in the tests are presented in Tab. I.
@@ IV-D4 Inference Time
This section quantifies the inference characteristics of the tested networks, including the number of parameters (# Parameters), multiply-accumulate operations (MACs), and inferencing speed. The unit of inference speed is frames per second (FPS). The number of classes was set to eight, and other configurations were consistent with those described in Sec. IV-A2. The inference speed results, averaged over 1000 runs, are presented in Tab. VII.
Methods | # Parameters | MACs | Speed |
/Million | /Billion | /FPS | |
UNet (MICCAI’15) [10] | 34.53 | 262.21 | 93 |
UNet++ (TMI’19) [11] | 9.16 | 139.68 | 100 |
PraNet (MICCAI’20) [12] | 30.34 | 25.65 | 82 |
HRNet (TPAMI’20) [29] | 63.60 | 65.80 | 78 |
nnUNet (NM’21) [13] | 30.60 | 232.19 | 83 |
CFANet (PR’23) [16] | 25.71 | 115.63 | 56 |
TransNetR (MIDL’24) [25] | 27.27 | 44.71 | 100 |
LiteSvANet (Ours) | 53.04 | 224.18 | 77 |
SvANet (Ours) | 177.64 | 312.76 | 55 |
Tab. VII illustrates that SvANet achieved a real-time analysis of medical images with 55 FPS. Notably, while SvANet consumed 312.76 billion MACs — two times more than CFANet — it performed at only 1 FPS lower than CFANet. This discrepancy highlights that MACs, as theoretical indicators of computational cost, do not fully capture the effects of hardware or software optimizations for inferencing. Despite the high computational load, SvANet’s performance remains well-suited for self-examination and clinical diagnostic applications.
Additionally, a streamlined version of SvANet, named LiteSvANet, was developed by omitting the fifth encoder stage while retaining the ASPP on the fourth encoder stage, reducing the parameter count by 70%. Subsequent tests, under identical conditions (described in Sec. IV-A2), demonstrated that LiteSvANet achieved a mDice of 70.88% and a sensitivity of 69.96% in the SpermHealth dataset, surpassing the performance of the second-best method, TransNetR, as shown in Tab. III. Moreover, LiteSvANet significantly enhanced the inference speed to 77 FPS, which, while 23% lower than the fastest models (UNet ++ and TransNetR), represents a considerable improvement over the standard SvANet model. For straightforward applications, implementing LiteSvANet is advantageous for examining small medical objects.
@@ IV-E Negative Case Studies
For tissue cell recognition in the TissueNet dataset, as shown in Fig. 9c, both TransNetR and SvANet effectively delineated cell boundaries and accurately labeled the cells and nuclei regions, closely resembling the ground truth. In contrast, other SOTA methods struggled to categorize cells and nuclei, leading to difficulties in differentiating cell boundaries and merging several cells. For sperm cell analysis, as presented by the final image in Fig. 9d, SvANet precisely located all sperm positions and effectively recognized the region of the short tail of an abnormal sperm. Conversely, tested methods like PraNet and CFANet struggled to differentiate the head and tail of the unhealthy sperm, as illustrated in the zoomed-in region of Fig. 9d. Moreover, UNet, UNet++, HRNet, nnUNet, and TransNetR misclassified an unhealthy sperm head as healthy, as indicated by a green subregion in Fig. 9d.